OpenStudy (anonymous):

At the park there is a pool shaped like a circle with radius 11yd . A ring-shaped path goes around the pool. Its width is 5yd .We are going to give a new layer of coating to the path. If one gallon of coating can cover 6yd2 , how many gallons of coating do we need? Note that coating comes only by the gallon, so the number of gallons must be a whole number. (Use the value 3.14 for π.)

OpenStudy (anonymous):

@KendrickLamar2014

OpenStudy (kendricklamar2014):

Note that coating comes only by the gallon, so the number of gallons must be a whole number. : With a 5' path around the pool the overall radius (including the path) 11 + 5 = 16 ft Path Area (pa) = Overall Area - Pool Area

OpenStudy (kendricklamar2014):

pa = (3.14*16^2) - (3.14*11^2) = pa = 803.84 - 379.94 = pa = 423.9 sq/yds

OpenStudy (kendricklamar2014):

To find the amount of coating needed: 423.9/6 = 70.65 = 71

OpenStudy (anonymous):

oh ok got it! I knew part of it but the the gallons part messed me up

OpenStudy (anonymous):

thank you!

OpenStudy (kendricklamar2014):

You're Welcome :)